Review Two small plastic spheres each have a mass of g and a charge of -53.0 nC. They are placed 2.7 cm apart (center to center) 2.0
▼ Part A What is the magnitude of the electric force on each sphere? Express your answer to two significant figures and include the appropriate units. HA F-1 Value Units Submit Request.Answen
Part B By what factor is the electric force on a sphere larger than its weight? Express your answer using two significant figures. TIn Submit st Answer
Answer #1

Answer :

1) We have F = k*q1*q2/r^2 = 9.0x10^9*(53.0x10^-9)^2/0.027^2 = 0.034679N

F = 3.5*10^-2 N

2) The weight = m*g = 0.002kg*9.8m/s^2 = 0.0196

So Fe/ Fg =0.034679/0.0196 = 1.796 = 1.8